40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Coating Ring-down Measurement Lab elog  Not logged in ELOG logo
Message ID: 323     Entry time: Tue Feb 28 14:20:50 2017
Author: Gabriele 
Type: General 
Category: Measurements 
Subject: New MIT sample #1 

2017-02-28

  • 2:19pm, installed in CR0, balanced, with 1/2" inserts facing down
    • It looks dusty and there is a fingerprint on the edge. I didn't clean it
  • 2:21pm, roughing pump on
  • 2:35pm, turbo pump on
  • Excitations:
    • Quiet time before excitation: 1172377933
      Excitation broadband: 1172377965
      Quiet time after excitation: 1172378027

    • Quiet time before excitation: 1172381657
      Excitation broadband: 1172381689
      Quiet time after excitation: 1172381751

    • Quiet time before excitation: 1172385381
      Excitation broadband: 1172385414
      Quiet time after excitation: 1172385476

    • Quiet time before excitation: 1172389106
      Excitation broadband: 1172389138
      Quiet time after excitation: 1172389200

    • Quiet time before excitation: 1172392830
      Excitation broadband: 1172392863
      Quiet time after excitation: 1172392925

    • Quiet time before excitation: 1172396555
      Excitation broadband: 1172396587
      Quiet time after excitation: 1172396649

 

2017-03-01

  • 9:36am, pumps stopped

2017-03-02

  • 9:00am. The laser beam was probably hitting one of the glued inserts. Checked that te disk is not touching, moved the ESD a bit lower
  • 9:04am, roughing pump on
  • 9:17am, turbo pump on
  • 1:50pm, pumps off
ELOG V3.1.3-